# Extended Fractional Chern Insulators Near Half Flux in Twisted Bilayer Graphene Above the Magic Angle

### v2

@@ V Discussion of possible mechanisms for extended FCI
S9 Vertical lines
In Fig. 3A, a dashed ellipse is drawn which encircles three sets of bands: one narrow red band near -343 meV, and two broader sets of bands between -330 and -320 meV. The narrow red band carries a Chern number of -4 and a net spin down, while the two broadened bands each carry a Chern number of 2 and have opposite spins. Thus, if the Fermi level is set within the gap on the left side of the ellipse, there is no net Chern number: $-4+2+2=0$. However, there is a net spin: the broadened band’s spins cancel each other, but there is no spin to cancel the narrow red band’s spin. This net property will hold even if the higher-energy bands are modified by interactions. All four gaps with this property are marked with Xs in panels A and B.
This situation is reminiscent of a quantum spin Hall insulator. We might expect, therefore, quantized longitudinal along with zero transverse resistance. In our longitudinal measurements, we tend to see vertical lines of minimal resistivity, often below our measurement floor of roughly 1 Ohm, however in many cases the minima is higher. Unfortunately, because of measurement noise and differences between neighboring contact pairs, our measurements are insufficient to make strong claims about the transport properties of these states.
Fig. S13 shows zoom-in regions just below half flux near $n/n_{s}=\pm 2$ for the high-quality fan in Fig. 1 and the good Hall pair from Fig. 2. At $n/n_{s}=-2$, $\rho_{xx}$ drops below our measurement floor of roughly 1 Ohm (panel A), and the Hall resistance is zero (panel C). At $n/n_{s}=2$, there is a vertical-line transition from low to high resistance rather than an isolated minimum (panel B), and the Hall resistance is not zero (panel D). Zero longitudinal and Hall resistance simultaneously would be some exotic superconducting state, but we do not claim to have such a state here. Rather, we expect our measurements are a result of inhomogeneity within the device.
Many of the other longitudinal contact pairs show minima in resistivity, some dropping below our measurement floor of 1 Ohm (Fig. S2). However, from one contact pair to the next, these minima vary by orders of magnitude. The top row of Fig. S14 shows $\rho_{xx}$ for all contact pairs near the four locations with vertical lines. The cuts are noisy and mostly unimpressive. Only contact pair 7 - 17 (the good pair with FCI) shows a clear zero resistance plateau at $n/n_{s}=-2$.
S10 How well quantized is the Hall resistance?
@@
S11 Discussion of other fractions
@@
S12 Modeling a domain wall
@@
S13 Spin dependent transport detail
As an example, the line cut along the white dotted line in Fig. 1 is shown in Fig. S27 panel A. To the left of $(0,4)$, the resistivity is peaks around 4 kilohms. To the right, between $(0,4)$ and $(2,0)$, it is only a hundred Ohms or lower. In panel B we show that these two types of transport correspond to states in the model that are entirely spin-polarized. Where the states are spin-up, we see low resistivity, and where they are spin-down we see high resistivity. Where we see a mixture of both, the resistivity is even higher. This behavior is followed in our measurements not just near $(0,\pm 4)$, but also in other regimes where the model reasonably matches the experiment. The major exception to this rule is near charge neutrality, where no substantial asymmetry is apparent.
Curiously, we only see FCI states in these low-resistivity (spin-up) regions. Circles in Fig. 3 panels A and B indicate regions where we observe ±$\pm 8/3$ FCI states. All of these regions are within blue Chern bands in panel A.
Spin-dependent resistivity has been observed in quantum Hall systems previously [58, 59, 60]. With Zeeman splitting, the lower-energy, spin-up states are filled first. Spin-down electrons then see a different magnetic environment: there is a substantial background of spin-up electrons already present. In high-quality WSe2, this leads to the minority-spin resistance being dramatically increased at low temperatures [60]. We observe the same behavior here, however we are not confident that the underlying explanation is the same. We may be able to reverse the direction of resistivity with back gate voltage. We wish to emphasize that we do not have positive evidence for the hypothesis of spin-dependent transport.
S14 Continuum Model and Quantum metric
@@
where $R_{\varphi}=\begin{pmatrix}\cos\varphi&-\sin\varphi\\ \sin\varphi&\cos\varphi\end{pmatrix}$ is the two-dimensional rotational matrix with the angle of $\varphi$ that determines the two principal axes of the strain matrix $S_{\epsilon}$, and $R_{\varphi}^{T}$ is the transpose of the matrix $R_{\varphi}$, $\nu=0.16$ is the Poisson ratio as used in Ref. [37], and $\epsilon_{\mathrm{uni}}$ and $\epsilon_{\mathrm{bi}}$ are the uniaxial strain and biaxial strain respectively. In principle, the two layers may have different strain matrices $S_{\epsilon}^{(t)}$ and $S_{\epsilon}^{(b)}$, where the superscript $t$ and $b$ refers to top and bottom layers respectively. However, in this work, we consider only heterostrain with $S_{\epsilon}=S_{\epsilon}^{(t)}-S_{\epsilon}^{(b)}$ and neglect the homostrain $\frac{1}{2}(S_{\epsilon}^{(t)}+S_{\epsilon}^{(b)})$, since the shape of moire unit cell, the narrow band structure, the associated band properties, etc. are more sensitive to heterostrain whose impact is estimated as the maximal value of $\epsilon/\theta$ with $\epsilon=max(|\epsilon_{uni}|,|\epsilon_{bi}|)$ is the larger one between the uniaxial strain and the biaxial strain. On the other hand, the impact of homostrain is estimated as $\sim\epsilon$, much smaller than the heterostrain, and therefore can be safely neglected for small twist angles.
@@
In Ref. [37], we estimated an uncertainty of roughly $\delta\nu=\pm 0.05$ on each van Hove point. Here, our best-fit result does satisfy this bound on each van Hove point.
We hesitate to place uncertainties on our best-fit parameters, because we believe that they will be dominated by systematics. The model is not perfect, and it is not straightforward to account for model uncertainty. The uncertainties that we do report are based on the width of the minima in Fig. S19, and are underestimates.
@@
Figure S1: Longitudinal fan with schematic. (A) Same data as in Fig. 1. (B) Black lines are ordinary integer quantum Hall. Blue vertical lines are potential QSH sites (Sec. S9). Green dashed lines are correlated Hofstadter ferromagnetic states (Sec. S5). Red dotted lines are symmetry-broken Chern insulating states (Sec. S5). Red dashed lines are the bending Landau levels discussed in Sec. IV. Gray dashed lines are Brown-Zak oscillations.
@@
Figure S5: Twist angle variation. Blue-labelled points are our contacts. The pitch between contacts is 3 $\mu$m. The channel is 1 $\mu$m wide. Vertical lines between them are Hall pairs, and horizontal lines are longitudinal pairs. The estimated $\theta$ values for each contact pair (Table S1) are laid over the lines, with uncertainties elided for brevity. Shaded regions indicate our guess of regions with different twist angles, based on nearby twist angles and the sharpness of features in the measurements.
@@
Figure S7: Hall plateaus. All line cuts at constant field of a Landau fan for a given contact pair plotted over each other with transparency. Quantized Hall plateaus will appear as solid horizontal lines. Contact pairs indicated in the top left corner. (A, B) There are no quantized Hall plateaus for these contact pairs, not even integers. (C) Contact pair 7 - 17 shows quantized plateaus at integers ±$\pm 8$ (not shown), ±$\pm 4$, -2, -8/3, -8/5, and -4/3, the latter two only near $n/n_{s}=3$. (D) Histogram of all Hall conductances. Interestingly, there are more -8/3 measurements than any integer.
@@
Figure S10: Detail of LL reset. Data from Fig. 1. Other contact pairs shown in Figs. S11 and S12. The green dashed line $(1,3)$ is not prominently visible in the resistivity, but lines up with the kink in the red dotted line below it. The red dotted lines indicate minima in $\rho_{\mathrm{XX}}$ and likely correspond to Landau level gaps. The black dotted lines are minima that follow the indicated paths
@@
Figure S19: Best fit to low-field van Hove points. (A-C) $\chi^{2}$ metric for the two indicated parameters (yellow is low, purple is high), using the best value of the not shown third dimension. (D-F) One dimensions $chi^{2}$ metric, using the best value of both not shown dimensions.
Figure S21: Raw Hall data. Unfiltered version of Fig. 2.
Figure S23: Highlighting values of Fractional Hall resistance In a similar manner to Fig. 2A of the main text, the longitudinal resistance is shown in grey-scale. We then highlight regions where the Hall resistance falls within 1% of the specified fractional value: A ±$\pm 8/3$ (reproduced from main text), B ±$\pm 8/5$, C ±$\pm 8/7$, D $4/3$, E $5/3$, and F 7/3. Our clustering algorithm has been applied to remove spurious points (see Sec. S3).
Figure S27: Possible spin-dependent transport. (A) Line cut from Fig. 1 along the dotted white line at 19.2 T. Intersections with the $(0,4)$, $(2,0)$, and $(4,-4)$ Hofstadter gaps are indicated with vertical dotted lines. (B) Density of states at $p/q=19/45$ ($B\approx 19.1$ T). Spin up and down states are indicated in blue and orange respectively, as well as with arrows in both panels.

### v3

@@ V Discussion of possible mechanisms for extended FCI
To show that the Landau levels emanating from half flux are two-fold degenerate (per spin per valley), the simplest way to see this is to look at the directly at the Hofstadter spectrum. For example, when crossing from the (4, -4) to the (0, 4) gap slightly above or below half flux, we cross a single Hofstadter subband (in the absence of Zeeman splitting). This corresponds to a total change in the Chern number of 8, so a single Landau level must carry an extra two-fold degeneracy.
To show this more concretely, we can start by considering a square lattice for simplicity. The effect of a magnetic field on the translation operators can be incorporated through a Peierls substitution
$T_{x}=e^{i(p_{x}+eA_{x})x/\hbar},\ T_{y}=e^{i(p_{y}+eA_{y})y/\hbar}$ (3)
where $A$ is the vector potential for some choice of gauge. For a given flux through the unit cell $\Phi/\Phi_{0}$, these operators will satisfy the commutation relation:
$T_{x}T_{y}=T_{y}T_{x}e^{i2\pi\Phi/\Phi_{0}}$ (4)
In the case where the flux per unit cell is a rational fraction $\Phi/\Phi_{0}=p/q$, for co-prime integers $p$ and $q$, this commutation relation can be used to show
$T_{x}^{q}T_{y}=T_{y}T_{x}^{q};$ (5)
commutativity is restored over the enlarged magnetic unit cell.
If we specifically consider the Landau gauge and consider a specific wavefunction $\psi_{n,\,p_{y}}$ that is an eigenstate of both the Hamiltonian and $T_{y}$. While this state is also an eigenstate of $T_{x}^{q}$, it is not an eigenstate of $T_{x}$. Therefore, there must exist some energetically degenerate state $T_{x}\psi_{n,\,p_{y}}$ with a different value of $p_{y}$. A state (or small cluster of states) that defines a low energy pocket of a band from which a sequence of Landau levels would form must be $q$-fold degenerate. In the case of half-flux, this means that a sequence of Landau levels has an extra two-fold degeneracy compared to a typical sequence emanating from zero field.
S8 Cluster filtering
The unfiltered data from Fig. 2 can be seen in Fig. S16. To filter the data, we begin by identifying points where the Hall resistance $R_{xy}$ falls within some window of our desired value (±$\pm 1\%$ of ±$\pm 3h/8e^{2}$ in this case). Binary dilation is applied to the windowed data such that nearest-neighbors within this masked array are now included. Each contiguous cluster of the dilated data set is then identified, and, if it does not contain above a threshold number of points, the original undilated points are omitted from the final filtered data set. Because of the dilation and generous threshold, this filtering primarily serves to remove isolated data points with no next-nearest-neighbor point within the window.
S9 How well quantized is the Hall resistance?
@@
S10 Discussion of other fractions
@@
S11 Modeling a domain wall
@@
S12 Continuum Model and Quantum metric
@@
where $R_{\varphi}=\begin{pmatrix}\cos\varphi&-\sin\varphi\\ \sin\varphi&\cos\varphi\end{pmatrix}$ is the two-dimensional rotational matrix with the angle of $\varphi$ that determines the two principal axes of the strain matrix $S_{\epsilon}$, and $R_{\varphi}^{T}$ is the transpose of the matrix $R_{\varphi}$, $\nu=0.16$ is the Poisson ratio as used in Ref. [45], and $\epsilon_{\mathrm{uni}}$ and $\epsilon_{\mathrm{bi}}$ are the uniaxial strain and biaxial strain respectively. In principle, the two layers may have different strain matrices $S_{\epsilon}^{(t)}$ and $S_{\epsilon}^{(b)}$, where the superscript $t$ and $b$ refers to top and bottom layers respectively. However, in this work, we consider only heterostrain with $S_{\epsilon}=S_{\epsilon}^{(t)}-S_{\epsilon}^{(b)}$ and neglect the homostrain $\frac{1}{2}(S_{\epsilon}^{(t)}+S_{\epsilon}^{(b)})$, since the shape of moire unit cell, the narrow band structure, the associated band properties, etc. are more sensitive to heterostrain whose impact is estimated as the maximal value of $\epsilon/\theta$ with $\epsilon=\mathrm{max}(|\epsilon_{uni}|,|\epsilon_{bi}|)$ is the larger one between the uniaxial strain and the biaxial strain. On the other hand, the impact of homostrain is estimated as $\sim\epsilon$, much smaller than the heterostrain, and therefore can be safely neglected for small twist angles.
@@
In Ref. [45], we estimated an uncertainty of roughly $\delta\nu=\pm 0.05$ on each van Hove point. Here, our best-fit result does satisfy this bound on each van Hove point. We hesitate to place uncertainties on our best-fit parameters, because we believe that they will be dominated by systematics. The model is not perfect, and it is not straightforward to account for model uncertainty. The uncertainties that we do report are based on the width of the minima in Fig. S25, and are underestimates.
The fits are only performed using low-field data. We find that at high field, the computed results look qualitatively similar for the amount of strain estimated in Ref. [45] and the amount fit here. See Fig. S24.
@@
S13 Spin dependent transport detail
As an example, a line cut at constant field is shown in Fig. S26 panel A. To the left of $(0,4)$, the resistivity is peaks around 4 kilohms. To the right, between $(0,4)$ and $(2,0)$, it is only a hundred Ohms or lower. In panel B we show that these two types of transport correspond to states in the model that are entirely spin-polarized. Where the states are spin-up, we see low resistivity, and where they are spin-down we see high resistivity. Where we see a mixture of both, the resistivity is even higher. This behavior is followed in our measurements not just near $(0,\pm 4)$, but also in other regimes where the model reasonably matches the experiment. The major exception to this rule is near charge neutrality, where no substantial asymmetry is apparent.
Curiously, we only see FCI states in these low-resistivity (spin-up) regions. Circled $\alpha$, $\beta$, and $\gamma$ in Fig. 3 panels A and B indicate regions where we observe ±$\pm 8/3$ FCI states. All of these regions are within blue Chern bands.
Spin-dependent resistivity has been observed in quantum Hall systems previously [71, 72, 73]. With Zeeman splitting, the lower-energy, spin-up states are filled first. Spin-down electrons then see a different magnetic environment: there is a substantial background of spin-up electrons already present. In high-quality WSe2, this leads to the minority-spin resistance being dramatically increased at low temperatures [73]. We observe the same behavior here, however we are not confident that the underlying explanation is the same. We may be able to reverse the direction of resistivity with back gate voltage. See Fig. S3 panels C and D from Ref. [44], where oscillations appear as a function of displacement field (y-axis). We wish to emphasize that we do not have positive evidence for the hypothesis of spin-dependent transport.
Figure S1: Twist angle variation. Blue-labelled points are our contacts. The pitch between contacts is 3 $\mu$m. The channel is 1 $\mu$m wide. Vertical segments between them are Hall pairs, and horizontal segments are longitudinal pairs. The estimated $\theta$ values for each contact pair (Table S1) are laid over the lines, with uncertainties elided for brevity. Shaded regions indicate our guess of regions with particularly inhomogeneous twist angles, based on nearby twist angles and the sharpness of features in the transport measurements.
Figure S3: Longitudinal fan with schematic. (A) Same data as in Fig. 1. (B) Black lines are ordinary integer quantum Hall. Blue vertical lines are potential QSH sites discussed in the main text. Green dashed lines are correlated Hofstadter ferromagnetic states (Sec. S6). Red dotted lines are symmetry-broken Chern insulating states (Sec. S6). Red dashed lines are the bending Landau levels discussed in Sec. IV. Gray dashed lines are Brown-Zak oscillations. (C) The solid black lines from B overlaid on the computation results as white dotted lines.
@@
Figure S8: Detail of LL reset. Data from Fig. 1. Other contact pairs shown in Figs. S9 and S10. The green dashed line $(1,3)$ is not prominently visible in the resistivity, but lines up with the kink in the red dotted line below it. The red dotted lines indicate minima in $\rho_{\mathrm{XX}}$ and likely correspond to Landau level gaps. The black dotted lines are minima that follow the indicated paths
@@
Figure S13: Hall plateaus. All line cuts at constant field of a Landau fan for a given contact pair plotted over each other with transparency. Quantized Hall plateaus will appear as solid horizontal lines. Contact pairs indicated in the top left corner. (A, B) There are no quantized Hall plateaus for these contact pairs, not even integers. (C) Contact pair 7 - 17 shows quantized plateaus at integers ±$\pm 8$ (not shown), ±$\pm 4$, -2, -8/3, -8/5, and -4/3, the latter two only near $n/n_{s}=3$. (D) Histogram of all Hall conductances from panel B. Interestingly, there are more -8/3 measurements than any integer.
@@
Figure S16: Raw Hall data. Unfiltered version of Fig. 2.
Figure S18: Highlighting values of Fractional Hall resistance In a similar manner to Fig. 2A of the main text, the longitudinal resistivity is shown in grey-scale. We then highlight regions where the Hall resistance falls within 1% of the specified fractional value: A ±$\pm 8/3$ (reproduced from main text), B ±$\pm 8/5$, C ±$\pm 8/7$, D $4/3$, E $5/3$, and F 7/3. Our clustering algorithm has been applied to remove spurious points (see Sec. S8). Note that the 1% cutoff leads to different sizes of ranges of resistances for each fraction.
@@
Figure S24: Computation comparison with different strain matrices. (Left panels) Spectra for the indicated strain parameters. (Right panels) Associated Wannier plots.
Figure S25: Best fit to low-field van Hove points. (A-C) $\chi^{2}$ metric for the two indicated parameters (yellow is low, purple is high), using the best value of the not shown third dimension. (D-F) One dimensions $\chi^{2}$ metric, using the best value of both not shown dimensions.
Figure S26: Possible spin-dependent transport. (A) Line cut from Fig. 1 at 19.2 T. Intersections with the $(0,4)$, $(2,0)$, and $(4,-4)$ Hofstadter gaps are indicated with vertical dotted lines. (B) Density of states at $p/q=19/45$ ($B\approx 19.1$ T). Spin up and down states are indicated in blue and orange respectively, as well as with arrows in both panels.